OpenStudy (anonymous):

cos2(theta) = 1/2, Solve for theta?

OpenStudy (anonymous):

\[\cos(2\Theta) = 2\cos^{2}\Theta -1\] so \[2\cos^{2}\Theta -1 = \frac{ 1 }{ 2 }\] can you solve it now?

OpenStudy (anonymous):

Add 1, divide by 2, then take the square root? Then we take inverse cos?

OpenStudy (anonymous):

\[\frac{ \sqrt{3} }{ 2 }\] When you solve at the end you get pi/6

OpenStudy (anonymous):

Think that's right thank you!

OpenStudy (anonymous):

Take the positive and negative roots so there are 3 more solutions.

OpenStudy (anonymous):

Thank you! I got it!

OpenStudy (anonymous):

niceee yw =))